5. Using 3-point formulas, complete the following table with estimates for f and f. f(x)| f(x) | S(x) 2.0 6.00 2.2 6.02 2
Answer #1

Svin - Three both to point before better formula allow us to and after the point evaluate the derinative. use to informationwhen x= 204 fin) = f(2-6) -412-2) 2x 0.2 = 6•12 - 6.02 : 2x 0.2 - 025 L ifan = f(2:2) -2of(2.0) + f(2-6) (0. 2² 6.02 - 2x 6.0ANSWER f(202) = 0.15. f (202) = 0.5. f (204) = 005 f(204) = 0025 f(2-6) = 0.25 f(2-6) = -005 CS Scanned with CamScanner